# Wasserstein Evolution : Evolutionary Optimization as Phase Transition

### v1

@@ Abstract
This paper establishes a novel connection between evolutionary computation and statistical physics by formalizing evolutionary optimization as a phase transition process. We introduce Wasserstein Evolution (WE), a principled optimization framework that implements the Wasserstein gradient flow of a free energy functional, mathematically bridging evolutionary dynamics with thermodynamics. WE directly translates the physical competition between potential gradient forces (exploitation) and entropic forces (exploration) into algorithmic dynamics, providing an adaptive, theoretically grounded mechanism for balancing exploration and exploitation. Experiments on challenging benchmark functions demonstrate that WE achieves competitive convergence performance while maintaining dramatically higher population diversity than classical methods (GA, DE, CMA-ES).This superior entropy preservation enables effective navigation of multi-modal landscapes without premature convergence, validating the physical interpretation of optimization as a disorder-to-order transition. Our work provides not only an effective optimization algorithm but also a new paradigm for understanding evolutionary computation through statistical physics.
@@ 1  Introduction
Inspired by this question, this study proposes Wasserstein Evolution (WE), a novel optimization framework that bridges evolutionary computation and statistical physics through the principles of Wasserstein gradient flow and free energy minimization. Experimental results demonstrate that, compared to classical evolutionary methods such as Genetic Algorithm (GA)[17], Differential Evolution (DE)[18], and Covariance Matrix Adaptation Evolution Strategy (CMA-ES)[19], WE achieves excellent convergence performance on a suite of benchmark functions including Rastrigin, Beale, Himmelblau, Six-Hump Camel Back, and Holder Table. Notably, WE maintains highest entropy upon convergence, enabling it to better capture multimodal solutions and avoid premature convergence. The dynamic evolution of population entropy further reflects a phase-transition-like progression from disorder to order, consistent with the theoretical underpinnings of the algorithm.The main contributions of this paper are summarized as follows:
@@
    We design the Wasserstein Evolution, which adaptively balances exploration and exploitation by translating the competition between potential gradient forces and entropic forces—central to physical systems—into an evolutionary optimization context.
@@
    We conduct extensive experiments on classical multimodal benchmark functions, systematically validating WE’s performance advantages—particularly its ability to sustain population diversity—and offering a new solution approach for complex optimization problems.
@@ 3  Evolutionary Optimization as Phase Transition
3.1  Wasserstein Gradient Flow
The Wasserstein gradient flow provides a rigorous geometric framework for describing the dynamical evolution of probability distributions. It generalizes the classical notion of gradient descent from Euclidean space to the space of probability measures endowed with the Wasserstein metric[29, 30, 31].
Given a functional $\mathcal{F}[\rho]$ defined on the space of probability densities $\rho(\mathbf{x})$, the Wasserstein gradient flow describes how $\rho$ evolves in time to minimize $\mathcal{F}$. The evolution is governed by the continuity equation:
$\frac{\partial\rho_{t}}{\partial t}=\nabla\cdot\left(\rho_{t}\nabla\frac{\delta\mathcal{F}[\rho_{t}]}{\delta\rho}\right),\qquad t\geq 0,$ (1)
where $\frac{\delta\mathcal{F}}{\delta\rho}$ denotes the first variation of the functional, and $\rho_{t}(\mathbf{x})=\rho(\mathbf{x},t)$ represents the probability density at time $t$.
In the context of evolutionary optimization, we consider the following free energy functional, which naturally incorporates both exploitation (fitness optimization) and exploration (entropy maximization):
$\mathcal{F}_{\beta}[\rho]=\underbrace{\mathbb{E}_{\mathbf{x}\sim\rho}[f(\mathbf{x})]}_{\text{Potential Energy (Exploitation)}}-\frac{1}{\beta}\underbrace{S[\rho]}_{\text{Entropy (Exploration)}},$ (2)
where $f(\mathbf{x})$ is the fitness function, $S[\rho]=-\int\rho(\mathbf{x})\log\rho(\mathbf{x})\,d\mathbf{x}$ is the Shannon entropy, and $\beta>0$ is the inverse temperature parameter controlling the trade-off between the two competing terms.
The Wasserstein gradient flow for this free energy functional leads to the following particle-level dynamics. Let $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}$ be $N$ particles approximating the distribution $\rho_{t}$. Their trajectories follow the mean-field equation:
$\frac{d\mathbf{x}_{i}}{dt}=-\nabla f(\mathbf{x}_{i})-\frac{1}{\beta}\nabla\log\rho_{t}(\mathbf{x}_{i}),\qquad i=1,\dots,N.$ (3)
The first term $\nabla f(\mathbf{x}_{i})$ represents the potential gradient force, which drives particles toward regions of higher fitness (exploitation). The second term $\frac{1}{\beta}\nabla\log\rho_{t}(\mathbf{x}_{i})$ represents the entropic force (or diversity pressure), which pushes particles away from high-density regions toward less explored areas (exploration).
The connection to physical phase transitions becomes evident through the temperature parameter $\beta$. At high temperature ($\beta\to 0$), the entropic term $\frac{1}{\beta}\nabla\log\rho$ dominates, resulting in a disordered phase characterized by broad exploration. As the system cools ($\beta$ increases), the potential gradient term becomes increasingly dominant, leading to an ordered phase focused on local exploitation. This continuous shift from exploration to exploitation mirrors the liquid-to-solid phase transition in physical systems.
The Wasserstein metric ensures that the flow follows the path of steepest descent with respect to the intrinsic geometry of probability distributions, making it a natural framework for describing population dynamics in optimization.
3.2  Fokker-Planck Equation
The mathematical formalism of Wasserstein gradient flow establishes a profound connection with the Fokker-Planck equation from statistical physics, revealing that these two seemingly distinct formulations describe the same underlying physical process from complementary perspectives. This connection provides crucial insights into the thermodynamic interpretation of evolutionary optimization[32].
For the specific free energy functional $\mathcal{F}_{\beta}[\rho]=\mathbb{E}_{\rho}[f]-\frac{1}{\beta}S[\rho]$, we compute its first functional derivative:
$\frac{\delta\mathcal{F}_{\beta}[\rho]}{\delta\rho}=f(\mathbf{x})+\frac{1}{\beta}(\log\rho(\mathbf{x})+1).$ (4)
Substituting this expression into the Wasserstein gradient flow equation yields:
$\frac{\partial\rho_{t}}{\partial t}$ $=\nabla\cdot\left[\rho_{t}\nabla\left(f(\mathbf{x})+\frac{1}{\beta}(\log\rho_{t}(\mathbf{x})+1)\right)\right]$
$=\nabla\cdot\left[\rho_{t}\nabla f(\mathbf{x})\right]+\frac{1}{\beta}\nabla\cdot\left[\rho_{t}\nabla\log\rho_{t}(\mathbf{x})\right]$
$=\nabla\cdot\left[\rho_{t}\nabla f(\mathbf{x})\right]+\frac{1}{\beta}\nabla^{2}\rho_{t}(\mathbf{x}),$ (5)
where the final equality follows from the identity $\rho_{t}\nabla\log\rho_{t}=\nabla\rho_{t}$. Equation (5) is precisely the Fokker-Planck (or Smoluchowski) equation:
@@ 3.2  Fokker-Planck Equation
with diffusion coefficient $D=\frac{1}{\beta}$.
This Fokker-Planck equation describes the probability density evolution for particles undergoing overdamped Langevin dynamics:
$d\mathbf{x}_{t}=-\nabla f(\mathbf{x}_{t})dt+\sqrt{2D}d\mathbf{W}_{t},$ (7)
where $\mathbf{W}_{t}$ is a standard Wiener process. The mathematical equivalence can be formally established through Itô’s formula: for any smooth test function $\phi(\mathbf{x})$, applying Itô’s lemma to $\phi(\mathbf{x}_{t})$ and taking expectations yields the Kolmogorov backward equation, whose dual is precisely the Fokker-Planck equation (6).
The inverse temperature parameter $\beta$ plays a dual role in this correspondence. In the Wasserstein formulation, $\beta^{-1}$ weights the relative strength of the entropic force $\frac{1}{\beta}\nabla\log\rho_{t}$ against the potential gradient force $\nabla f$. In the Fokker-Planck formulation, $\beta^{-1}=D$ represents the diffusion coefficient controlling the intensity of stochastic fluctuations.
The stationary solution of the Fokker-Planck equation (6) is the Gibbs distribution:
$\rho_{\infty}(\mathbf{x})=\frac{1}{Z}e^{-\beta f(\mathbf{x})},\qquad Z=\int_{\Omega}e^{-\beta f(\mathbf{x})}d\mathbf{x},$ (8)
which minimizes the free energy $\mathcal{F}_{\beta}[\rho]$. This establishes a direct link between the long-time behavior of the stochastic dynamics and the equilibrium statistical mechanics of the system.
The convergence properties can be analyzed through the free energy dissipation. Along solutions of the Fokker-Planck equation, the free energy decreases monotonically:
$\frac{d}{dt}\mathcal{F}_{\beta}[\rho_{t}]$ $=\int_{\Omega}\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\frac{\partial\rho_{t}}{\partial t}d\mathbf{x}$
@@
$=\int_{\Omega}\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\nabla\cdot\left[\rho_{t}\nabla\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\right]d\mathbf{x}$
$=-\int_{\Omega}\rho_{t}\left\|\nabla\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\right\|^{2}d\mathbf{x}\leq 0,$ (9)
where the final equality follows from integration by parts. This dissipation inequality quantifies the rate at which the system approaches equilibrium and provides a natural convergence criterion for optimization algorithms.
This mathematical equivalence reveals that evolutionary optimization can be viewed through two complementary lenses: as a deterministic gradient flow in the Wasserstein geometry of probability distributions, or as a stochastic process described by the Fokker-Planck equation. The former emphasizes the collective behavior of the population as a whole, while the latter focuses on the individual particle dynamics. Both perspectives converge to the same physical reality: the minimization of free energy through the interplay of deterministic forces (exploitation) and stochastic fluctuations (exploration), with the inverse temperature $\beta$ governing the transition from disordered exploration to ordered exploitation.
3.3  Free Energy Minimization
The principle of free energy minimization provides a unified perspective on the convergence behavior of both physical systems and optimization algorithms[33]. In statistical physics, the equilibrium state of a system at temperature $T=1/\beta$ is given by the Gibbs distribution that minimizes the Helmholtz free energy:
$\rho_{\text{eq}}(\mathbf{x})=\frac{1}{Z}e^{-\beta f(\mathbf{x})},\qquad Z=\int e^{-\beta f(\mathbf{x})}d\mathbf{x},$ (10)
where $Z$ is the partition function. This distribution minimizes the functional:
$\mathcal{F}_{\beta}[\rho]=\mathbb{E}_{\rho}[f]-\frac{1}{\beta}S[\rho].$ (11)
The minimization process exhibits distinctive characteristics at different temperature regimes:
  • High temperature limit ($\beta\to 0$): The entropy term dominates, $\rho_{\text{eq}}(\mathbf{x})$ approaches a uniform distribution, maximizing exploration.
  • Low temperature limit ($\beta\to\infty$): The energy term dominates, $\rho_{\text{eq}}(\mathbf{x})$ concentrates around global minima of $f(\mathbf{x})$, maximizing exploitation.
  • Finite temperature: The distribution balances between exploration (high entropy) and exploitation (low energy).
The dynamical approach to equilibrium follows the Wasserstein gradient flow described in the previous section. This flow can be interpreted as a gradient descent in the space of probability distributions:
$\frac{d}{dt}\mathcal{F}_{\beta}[\rho_{t}]=-\int\rho_{t}(\mathbf{x})\left\|\nabla\frac{\delta\mathcal{F}_{\beta}[\rho_{t}]}{\delta\rho}\right\|^{2}d\mathbf{x}\leq 0,$ (12)
which guarantees monotonic decrease of free energy over time.
In evolutionary computation, the population distribution $\rho_{t}(\mathbf{x})$ evolves to minimize a similar objective. The fitness function $f(\mathbf{x})$ corresponds to the energy landscape, while population diversity corresponds to entropy. The selection pressure parameter in evolutionary algorithms plays a role analogous to the inverse temperature $\beta$. As the algorithm progresses, increasing selection pressure corresponds to decreasing temperature, driving the population from a disordered, exploratory phase to an ordered, exploitative phase.
This formal correspondence suggests that evolutionary optimization can be viewed as a non-equilibrium thermodynamic process, where the population undergoes a phase transition from high-entropy exploration to low-entropy exploitation. The competition between selection pressure (driving convergence) and mutation/crossover operators (maintaining diversity) parallels the competition between energy minimization and entropy maximization in physical systems.
The mathematical framework of free energy minimization provides a principled way to design optimization algorithms that naturally balance exploration and exploitation through temperature scheduling, much like simulated annealing but operating at the population distribution level rather than individual solution level.
@@ 3.4  Wasserstein Evolution
3:  Compute current inverse temperature: $\beta_{k}=\beta_{0}+(\beta_{\text{final}}-\beta_{0})\cdot(k/K)^{2}$
4:  Estimate density: $\hat{\rho}_{k}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}K_{h}(\mathbf{x}-\mathbf{x}_{i}^{(k)})$
@@
6:   Compute fitness gradient: $\nabla f(\mathbf{x}_{i}^{(k)})\approx\frac{f(\mathbf{x}_{i}^{(k)}+\epsilon)-f(\mathbf{x}_{i}^{(k)}-\epsilon)}{2\epsilon}$
7:   Compute entropic force: $\nabla\log\hat{\rho}_{k}(\mathbf{x}_{i}^{(k)})\approx\frac{\nabla\hat{\rho}_{k}(\mathbf{x}_{i}^{(k)})}{\hat{\rho}_{k}(\mathbf{x}_{i}^{(k)})+\delta}$
@@
9:   Apply boundary constraints: $\mathbf{x}_{i}^{(k+1)}\leftarrow\text{clip}(\mathbf{x}_{i}^{(k+1)},\mathbf{l},\mathbf{u})$
10:  end for
11:  Compute population statistics: entropy $S_{k}$, diversity $D_{k}$, best fitness $f_{k}^{*}$
12:end for
13:return Best solution $\mathbf{x}^{*}$ and corresponding fitness $f(\mathbf{x}^{*})$
The algorithm consists of three key components:
1. Density Estimation: At each iteration, we estimate the population density $\hat{\rho}_{k}(\mathbf{x})$ using kernel density estimation (KDE):
$\hat{\rho}_{k}(\mathbf{x})=\frac{1}{Nh^{d}}\sum_{i=1}^{N}K\left(\frac{\mathbf{x}-\mathbf{x}_{i}^{(k)}}{h}\right),$ (13)
where $K$ is a Gaussian kernel and $h$ is the bandwidth parameter. The gradient $\nabla\log\hat{\rho}_{k}(\mathbf{x})$ is then computed numerically using finite differences.
2. Temperature Scheduling: The inverse temperature $\beta_{k}$ follows an annealing schedule:
$\beta_{k}=\beta_{0}+(\beta_{\text{final}}-\beta_{0})\cdot\left(\frac{k}{K}\right)^{2},$ (14)
where $\beta_{0}$ is the initial inverse temperature (high exploration) and $\beta_{\text{final}}$ is the final inverse temperature (high exploitation). The quadratic scheduling ensures a smooth transition from exploration to exploitation, mimicking the cooling process in physical phase transitions.
3. Force Composition: Each particle experiences two competing forces:
$\mathbf{F}_{\text{total}}(\mathbf{x})=-\nabla f(\mathbf{x})-\frac{1}{\beta_{k}}\nabla\log\hat{\rho}_{k}(\mathbf{x}).$ (15)
The relative strength of these forces evolves over time, with the entropic force ($\frac{1}{\beta_{k}}\nabla\log\hat{\rho}_{k}$) dominating in early iterations (high temperature phase) and the potential gradient force ($\nabla f$) dominating in later iterations (low temperature phase).
Theoretical Properties:
Theorem 1 (Free Energy Decrease).
Under mild regularity conditions on $f$ and with appropriate learning rate $\eta_{k}$, the population distribution $\rho_{k}$ in WE satisfies:
$\mathcal{F}_{\beta_{k}}[\rho_{k+1}]\leq\mathcal{F}_{\beta_{k}}[\rho_{k}]-\frac{\eta_{k}}{2}\mathbb{E}_{\rho_{k}}\left[\|\mathbf{F}_{\text{total}}(\mathbf{x})\|^{2}\right]+O(\eta_{k}^{2}).$ (16)
Theorem 2 (Phase Transition Characterization).
Define the order parameter $m_{k}=\mathbb{E}_{\rho_{k}}[\|\mathbf{x}-\boldsymbol{\mu}_{k}\|^{2}]$ where $\boldsymbol{\mu}_{k}=\mathbb{E}_{\rho_{k}}[\mathbf{x}]$. As $\beta_{k}$ increases, WE exhibits:
  • Disordered phase ($\beta_{k}<\beta_{c}$): $m_{k}$ remains large, population entropy high
  • Critical point ($\beta_{k}\approx\beta_{c}$): Rapid decrease in $m_{k}$, entropy drops sharply
  • Ordered phase ($\beta_{k}>\beta_{c}$): $m_{k}$ small, population concentrated near optima

### v2

@@ Abstract
This paper introduces a novel framework linking evolutionary computation to statistical physics by formulating optimization as a statistical phase transition. We propose Wasserstein Evolution (WE), an algorithm based on the Wasserstein gradient flow of a free energy functional, translating the physical competition between exploitative potential gradient forces and explorative entropic forces into an adaptive search mechanism. Theoretical analysis confirms WE’s convergence to the Boltzmann distribution, ensuring a principled performance-diversity balance.Experiments compare WE against five established algorithms, including GA, DE, CMA-ES, JADE, and SaDE, on benchmark and physical potential functions. Results demonstrate that WE consistently achieves the lowest free energy and maintains the highest entropy, excelling in both solution quality and diversity preservation. Additional invariance tests using transformed Schwefel 2.22 functions verify that WE possesses translation, scale, and rotation invariance, proving its robustness and intrinsic alignment with the problem’s geometry rather than its coordinate representation.Thus, this work delivers not only an effective and robust optimizer but also a new theoretical paradigm for understanding population-based search through statistical physics, viewing convergence as a disorder-to-order phase transition and opening new avenues for designing intelligent optimization methods.
@@ 1  Introduction
Inspired by this question, this study proposes Wasserstein Evolution (WE), a novel optimization framework that bridges evolutionary computation and statistical physics through the principles of Wasserstein gradient flow and free energy minimization. Experimental results demonstrate that, compared to classical evolutionary methods such as Genetic Algorithm (GA)[17], Differential Evolution (DE)[18], Covariance Matrix Adaptation Evolution Strategy (CMA-ES)[19], Adaptive Differential Evolution with Optional External Archive (JADE)[20] and Self-adaptive Differential Evolution (SaDE)[21], WE achieves excellent convergence performance on a suite of benchmark functions including Rastrigin, Beale, Himmelblau, Six-Hump Camel Back, and Holder Table, as well as on physical potential functions such as the 2D Periodic, Double-Well, Tokamak, Multipole, and Optical Lattice potentials.Notably, WE achieves the lowest free energy while maintaining the highest entropy, enabling it to better capture multimodal solutions and avoid premature convergence.Theoretical analysis proves that WE converges to the Boltzmann distribution, a finding that aligns with our experimental observations. Furthermore, systematic tests on the Schwefel 2.22 function and its transformed variants verify that WE possesses translation, scale, and rotation invariance.The main contributions of this paper are summarized as follows:
@@
    We design the Wasserstein Evolution, which adaptively balances exploration and exploitation by translating the competition between potential gradient forces and entropic forces, central to physical systems, into an evolutionary optimization context.
@@
    We conduct extensive experiments on classical multimodal benchmark functions and real-world-inspired physical potential functions, systematically validating WE’s performance advantages.Specifically, WE consistently achieves the lowest free energy and highest entropy compared to established algorithms, including GA, DE, CMA-ES, JADE, and SaDE, demonstrating superior solution quality and sustained population diversity.
@@
    We demonstrate that WE possesses translation, scale, and rotation invariance through systematic tests on transformed variants of the Schwefel 2.22 function, confirming its robustness and that its dynamics are intrinsically tied to the geometry of the fitness landscape.
  5. 5.
    We provide a rigorous convergence analysis, proving that WE theoretically converges to the Boltzmann distribution—a result further corroborated by our experimental observations.
  6. 6.
@@ 3  Evolutionary Optimization as Phase Transition
3.1  Physical Background
The Wasserstein Evolution (WE) is grounded in three interconnected mathematical frameworks from statistical physics: Wasserstein gradient flow, the Fokker-Planck equation, and free energy minimization. Together, these provide a unified physical interpretation of evolutionary optimization as a phase transition process.
The Wasserstein gradient flow generalizes gradient descent to the space of probability measures. For a functional $\mathcal{F}[\rho]$ defined on probability distributions, its gradient flow follows:
$\frac{\partial\rho_{t}}{\partial t}=\nabla\cdot\left(\rho_{t}\nabla\frac{\delta\mathcal{F}[\rho_{t}]}{\delta\rho}\right),$ (1)
where $\frac{\delta\mathcal{F}}{\delta\rho}$ is the functional derivative and $\rho_{t}(\mathbf{x})$ is the probability density at time $t$. This establishes steepest descent in the Wasserstein geometry of probability spaces.
The Fokker-Planck equation describes probability density evolution for stochastic systems:
@@ 3.1  Physical Background
which governs the overdamped Langevin dynamics $d\mathbf{x}_{t}=-\nabla f(\mathbf{x}_{t})dt+\sqrt{2D}d\mathbf{W}_{t}$. Here, $f(\mathbf{x})$ represents the energy landscape and $D$ is the diffusion coefficient.
Free energy minimization characterizes equilibrium thermodynamics through the functional:
$\mathcal{F}_{\beta}[\rho]=\mathbb{E}_{\rho}[f(\mathbf{x})]-\frac{1}{\beta}S[\rho],$ (3)
where $S[\rho]$ is Shannon entropy and $\beta$ is inverse temperature. The minimizing distribution is the Gibbs equilibrium $\rho_{\text{eq}}(\mathbf{x})=Z^{-1}e^{-\beta f(\mathbf{x})}$.
These frameworks unify when we consider the specific free energy $\mathcal{F}_{\beta}[\rho]$. Its Wasserstein gradient flow becomes:
@@
$\frac{\partial\rho_{t}}{\partial t}=\nabla\cdot\left[\rho_{t}\nabla f(\mathbf{x})\right]+\frac{1}{\beta}\nabla^{2}\rho_{t},$ (4)
which is exactly the Fokker-Planck equation with $D=1/\beta$. This reveals that free energy minimization, implemented via Wasserstein gradient flow, produces the stochastic dynamics of the Fokker-Planck system. Consequently, WE can be viewed as a thermodynamic process where a population evolves to minimize free energy, undergoing a controlled phase transition from high-temperature, high-entropy exploration to low-temperature, low-entropy exploitation through inverse temperature annealing.
@@ 3.2  Wasserstein Evolution
3:  Compute current inverse temperature: $\beta_{k}$
4:  Estimate density: $\hat{\rho}_{k}(\mathbf{x})$
@@
6:   Compute fitness gradient: $\nabla f(\mathbf{x}_{i}^{(k)})$
7:   Compute entropic force: $\nabla\log\hat{\rho}_{k}(\mathbf{x}_{i}^{(k)})$
@@
9:  end for
10:  Return free energy $\mathcal{F}$, entropy $S$, best fitness $f^{*}$, best solution $\mathbf{x}^{*}$
11:end for
12:return Best solution $\mathbf{x}^{*}$ and corresponding fitness $f(\mathbf{x}^{*})$
The WE comprises three key components: (1) Density Estimation ($\hat{\rho}_{k}$), which approximates the population distribution at each iteration; (2) Temperature Scheduling ($\beta_{k}$), which controls the exploration-exploitation trade-off through an annealing scheme; and (3) Force Composition ($\nabla f-\frac{1}{\beta_{k}}\nabla\log\hat{\rho}_{k}$), which combines the gradient of the objective function with the entropic term to drive the population dynamics.
@@
3.3  Convergence Analysis of the Wasserstein Evolution
We provide a comprehensive convergence analysis of the Wasserstein Evolution (WE). The algorithm operates on a search space $\mathcal{X}\subseteq\mathbb{R}^{d}$ with objective function $f:\mathcal{X}\to\mathbb{R}$. We consider the minimization problem where we aim to find $\min_{x\in\mathcal{X}}f(x)$. At each iteration $t$, the algorithm maintains $N$ particles $X_{t}^{(1)},X_{t}^{(2)},\dots,X_{t}^{(N)}$ that evolve according to deterministic dynamics.
The core update equation for each particle is given by:
$\frac{dX^{(i)}}{dt}=-\nabla f(X^{(i)})-\frac{1}{\beta_{t}}\nabla\log\rho_{t}(X^{(i)}),$ (5)
where $\rho_{t}(x)$ represents the probability density of the particle distribution at time $t$, and $\beta_{t}>0$ is the inverse temperature parameter that varies with time. The first term $\nabla f(X^{(i)})$ drives particles toward regions of lower function values (for minimization problems), while the second term $\frac{1}{\beta_{t}}\nabla\log\rho_{t}(X^{(i)})$ acts as a repulsive entropic force that prevents particle concentration and maintains diversity.
Rather than analyzing individual particle trajectories, we consider the evolution of the entire probability distribution $\rho_{t}(x)$. The empirical distribution formed by the particles approximates this continuous density. The relationship between particle motion and distribution evolution is governed by the continuity equation, which states that the change in probability density equals the negative divergence of the probability flux. Substituting the velocity field from equation (5) yields the partial differential equation:
$\frac{\partial\rho_{t}(x)}{\partial t}=-\nabla\cdot\left[\rho_{t}(x)\left(-\nabla f(x)-\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)\right)\right].$ (6)
This equation describes how the probability distribution evolves over time in response to the combined forces of potential gradient and entropic repulsion.
To analyze convergence, we introduce the free energy functional defined as:
$\mathcal{F}_{\beta_{t}}[\rho]=\mathbb{E}_{x\sim\rho}[f(x)]-\frac{1}{\beta_{t}}S[\rho],$ (7)
where $S[\rho]=-\int\rho(x)\log\rho(x)dx$ is the (negative) Shannon entropy. The first term $\mathbb{E}_{x\sim\rho}[f(x)]=\int f(x)\rho(x)dx$ represents the average objective function value, while the second term $\frac{1}{\beta_{t}}S[\rho]$ represents the entropic contribution. Note that since $S[\rho]$ is negative, $S[\rho]/\beta_{t}$ is positive, which encourages entropy maximization. The free energy functional serves as a Lyapunov function for the dynamics.
We now compute the time derivative of the free energy along the solution trajectory of equation (6). First, we compute the functional derivative:
$\frac{\delta\mathcal{F}_{\beta_{t}}[\rho]}{\delta\rho(x)}=f(x)+\frac{1}{\beta_{t}}(\log\rho(x)+1).$ (8)
The time derivative is then:
$\frac{d}{dt}\mathcal{F}_{\beta_{t}}[\rho_{t}]=\int_{\mathcal{X}}\frac{\delta\mathcal{F}_{\beta_{t}}[\rho_{t}]}{\delta\rho(x)}\frac{\partial\rho_{t}(x)}{\partial t}dx.$ (9)
Substituting equation (6) gives:
$\frac{d}{dt}\mathcal{F}_{\beta_{t}}[\rho_{t}]$ $=\int_{\mathcal{X}}\left[f(x)+\frac{1}{\beta_{t}}(\log\rho_{t}(x)+1)\right]$ (10)
$\qquad\times\left(-\nabla\cdot\left[\rho_{t}(x)\left(-\nabla f(x)-\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)\right)\right]\right)dx.$ (11)
Applying integration by parts with the assumption that boundary terms vanish (which holds for compact domains or rapidly decaying distributions), we obtain:
$\frac{d}{dt}\mathcal{F}_{\beta_{t}}[\rho_{t}]$ $=\int_{\mathcal{X}}\nabla\left[f(x)+\frac{1}{\beta_{t}}(\log\rho_{t}(x)+1)\right]$ (12)
$\qquad\cdot\left[\rho_{t}(x)\left(-\nabla f(x)-\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)\right)\right]dx.$ (13)
Evaluating the gradient inside the integral yields:
$\frac{d}{dt}\mathcal{F}_{\beta_{t}}[\rho_{t}]$ $=\int_{\mathcal{X}}\rho_{t}(x)\left(\nabla f(x)+\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)\right)$ (14)
$\qquad\cdot\left(-\nabla f(x)-\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)\right)dx.$ (15)
This simplifies to:
$\frac{d}{dt}\mathcal{F}_{\beta_{t}}[\rho_{t}]=-\int_{\mathcal{X}}\rho_{t}(x)\left\|\nabla f(x)+\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)\right\|^{2}dx\leq 0.$ (16)
Equation (16) shows that the free energy is non-increasing along the trajectory, establishing $\mathcal{F}_{\beta_{t}}[\rho_{t}]$ as a valid Lyapunov function. The dissipation rate is zero if and only if $\nabla f(x)+\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)=0$ for all $x$ in the support of $\rho_{t}$.
At equilibrium, when $\frac{\partial\rho_{t}}{\partial t}=0$, equation (6) implies that the probability flux vanishes, leading to:
$\rho_{t}(x)\left(-\nabla f(x)-\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)\right)=0.$ (17)
Assuming $\rho_{t}(x)>0$ almost everywhere, this reduces to:
$\nabla f(x)-\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)=0,$ (18)
which is equivalent to $\nabla f(x)+\frac{1}{\beta_{t}}\nabla\log\rho_{t}(x)=0$, consistent with the condition for zero dissipation in equation (16).
Rearranging equation (18) gives $\nabla\log\rho_{t}(x)=-\beta_{t}\nabla f(x)$. Integrating both sides with respect to $x$ yields:
$\log\rho_{t}(x)=-\beta_{t}f(x)+C,$ (19)
where $C$ is an integration constant. Exponentiating produces:
$\rho_{t}(x)=e^{C}\cdot e^{-\beta_{t}f(x)}.$ (20)
Normalizing to obtain a valid probability density function gives the equilibrium distribution:
$\rho_{\text{eq}}(x)=\frac{1}{Z_{\beta_{t}}}e^{-\beta_{t}f(x)},$ (21)
where $Z_{\beta_{t}}=\int_{\mathcal{X}}e^{-\beta_{t}f(x)}dx$ is the partition function ensuring $\int\rho_{\text{eq}}(x)dx=1$.
This equilibrium distribution is precisely the Boltzmann distribution with energy function $f(x)$ (for minimization). The parameter $\beta_{t}$ controls the concentration of the distribution: when $\beta_{t}$ is small, the distribution is diffuse and exploratory; when $\beta_{t}$ is large, the distribution concentrates on regions where $f(x)$ is minimized.
To establish convergence to this equilibrium, we examine the behavior of the free energy functional at the equilibrium distribution. Substituting equation (21) into equation (7) gives:
$\mathcal{F}_{\beta_{t}}[\rho_{\text{eq}}]=\frac{1}{Z_{\beta_{t}}}\int f(x)e^{-\beta_{t}f(x)}dx-\frac{1}{\beta_{t}}S[\rho_{\text{eq}}].$ (22)
Using the properties of the Boltzmann distribution, we can show that for the specific form $\rho_{\text{eq}}(x)=e^{-\beta_{t}f(x)}/Z_{\beta_{t}}$, the entropy is given by:
$S[\rho_{\text{eq}}]=\log Z_{\beta_{t}}+\beta_{t}\mathbb{E}_{\rho_{\text{eq}}}[f(x)].$ (23)
Substituting equation (23) into the free energy expression yields:
$\mathcal{F}_{\beta_{t}}[\rho_{\text{eq}}]=-\frac{1}{\beta_{t}}\log Z_{\beta_{t}}.$ (24)
This represents the minimum possible value of the free energy functional for fixed $\beta_{t}$. Indeed, for any distribution $\rho$, we have:
$\mathcal{F}_{\beta_{t}}[\rho]-\mathcal{F}_{\beta_{t}}[\rho_{\text{eq}}]=\frac{1}{\beta_{t}}D_{\text{KL}}(\rho\|\rho_{\text{eq}})\geq 0,$ (25)
where $D_{\text{KL}}$ is the Kullback-Leibler divergence, which is non-negative.
The convergence argument proceeds as follows. Equation (16) shows that the free energy decreases monotonically along the trajectory. Combined with the lower bound in equation (25), this ensures convergence to the equilibrium distribution. Moreover, as $\beta_{t}$ increases over time (a process called annealing), the equilibrium distribution becomes increasingly concentrated on the global minimum of $f(x)$.
In the limit $\beta_{t}\to\infty$, the Boltzmann distribution converges weakly to a Dirac delta distribution centered on the global minimum:
$\lim_{\beta_{t}\to\infty}\rho_{\text{eq}}(x)=\delta(x-x^{*}),$ (26)
where $x^{*}=\arg\min_{x\in\mathcal{X}}f(x)$, assuming the minimum is unique. For multiple global minima, the distribution converges to a mixture of Dirac deltas at each global minimum.
The mathematical rigor of this convergence is supported by the fact that the dynamics constitute a gradient flow in the space of probability measures equipped with the Wasserstein metric. The free energy functional is geodesically convex under appropriate conditions on $f(x)$, ensuring exponential convergence to equilibrium. The discrete-time implementation with finite particle count introduces approximation errors, but these can be controlled through the choice of kernel bandwidth in density estimation and through the learning rate in the discretization of equation (5).
The WE thus provides a principled approach to global optimization that combines gradient-based local search with entropy-driven exploration. The theoretical guarantee of convergence to the Boltzmann distribution ensures that, with appropriate annealing of $\beta_{t}$, the algorithm will eventually concentrate its search effort on the globally optimal regions of the search space, while maintaining diversity through the entropy term to escape local optima during the initial phases of optimization.